# Towards Atomic MIMO Receivers

### v2

@@ III-A Architecture and Principles of Atomic MIMO Receiver
Due to the presence of reference signal ${\mathbf{b}}$, (18) is called a biased PR model as opposed to the typical PR model without this reference signal.
III-B Atomic-MIMO Detection Problems
The problem of atomic-MIMO detection is to estimate the multi-user data vector, ${\mathbf{s}}$, from the received signal, ${\mathbf{z}}$, in (18). Two problem formulations are provided below.
III-B1 Least-Square Criterion
@@
III-B2 Maximum-Likelihood Criterion
@@
Although problems (19) and (21) can be optimally solved by an exhaustive search, the complexity is prohibitive when the users are many and/or the constellation size is large. To this end, two low-complexity algorithms are designed in the next section.
IV Atomic-MIMO Detection Algorithms
In this section, we present two atomic-MIMO detection algorithms to solve the biased PR problems formulated in the preceding section. Their performance and complexity are subsequently analyzed.
IV-A LS Phase Retrieval Algorithm
@@
One issue in the implementation of the biased GS algorithm is its sensitivity to the choice of initial point. Without a proper initial estimate, Algorithm 1 may converge to an undesirable local optimum. To address this issue, the computationally efficient spectral method has been developed to yield good initialization [31]. Specifically, to adopt the spectral method into our model, we first construct an augmented observation matrix $\bar{{\mathbf{A}}}=[{\mathbf{A}}^{H},{\mathbf{b}}]^{H}$ as presented in Step 1. This matrix allows us to treat the user signal and reference signal together: $\bar{{\mathbf{s}}}=[{\mathbf{s}}^{H},1]^{H}$. Thereafter the observed amplitude vector can be rewritten in a reference-free form: ${\mathbf{z}}=|{\mathbf{A}}^{H}{\mathbf{s}}+{\mathbf{b}}+{\mathbf{w}}|=|\bar{{\mathbf{A}}}^{H}\bar{{\mathbf{s}}}+{\mathbf{w}}|$, whose initialization can be performed using the typical spectral method. To elaborate, define the weighted covariance matrix as ${\mathbf{M}}=\sum_{n=1}^{N}z_{n}\bar{{\mathbf{a}}}_{n}\bar{{\mathbf{a}}}_{n}^{H}$, where $\bar{{\mathbf{a}}}_{n}$ denotes the $n$-th column of $\bar{{\mathbf{A}}}$. The direction of the initial value $\bar{{\mathbf{s}}}^{0}$ of $\bar{{\mathbf{s}}}$ is estimated as the principal eigenvector ${\mathbf{v}}$ of ${\mathbf{M}}$. In addition, the magnitude $\bar{r}$ of $\bar{{\mathbf{s}}}^{0}$ is obtained by approaching $|{\mathbf{A}}^{H}\bar{{\mathbf{s}}}^{0}|=|{\mathbf{A}}^{H}\bar{r}{\mathbf{v}}|$ to ${\mathbf{z}}$:
@@
IV-B ML Phase Retrieval Algorithm
@@
where ${\mathbf{s}}^{t-1}$ denotes the previous estimate of ${\mathbf{s}}$. The EM algorithm proceeds iteratively by two steps to maximize the loglikelihood. They are the Expectation step (E-step) to derive the surrogate function $Q({\mathbf{s}}|{\mathbf{s}}^{t-1})$ by calculating the expectation in (IV-B), and the Maximization step (M-step) to update ${\mathbf{s}}^{t}$ by maximizing $Q({\mathbf{s}}|{\mathbf{s}}^{t-1})$, i.e., ${\mathbf{s}}^{t}=\arg\max_{{\mathbf{s}}}Q({\mathbf{s}}|{\mathbf{s}}^{t-1})$. These two steps are elaborated below.
  • E-step: In (IV-B), the likelihood function $p(z_{n},\theta_{n};{\mathbf{s}})$ stands for the ML estimator to ${\mathbf{s}}$ when both the amplitude $z_{n}$ and phase $\theta_{n}$ are accessible. Clearly, $p(z_{n},\theta_{n};{\mathbf{s}})$ is a Gaussian distribution
@@
    Moreover, to avoid convergence to a poor local optimum, the spectral method based on the augmented matrix $\bar{{\mathbf{A}}}$ as used in Section IV-A is carried out to obtain an initial estimate ${\mathbf{s}}^{0}$. In the end, after performing the initialization and EM-based iteration, the projection operation $\hat{s}_{k}=\arg\min_{s\in\mathcal{S}}\|\tilde{s}_{k}-s\|$ is applied to complete the constellation demapping.
  • IV-C Comparison between biased GS and EM-GS algorithms
@@
    IV-D Performance Analysis
    IV-D1 CRLB Analysis
@@
    where ${\mathbf{I}}$ denotes the Fisher information matrix, the $(p,q)$-th entry of which is determined by $I_{pq}=-\sum_{n=1}^{N}\mathsf{E}\left(\frac{\partial^{2}}{\partial s_{p}^{*}\partial s_{q}}\log p(z_{n};{\mathbf{s}})\right)$. An explicit expression of ${\mathbf{I}}$ is provided by the following Lemma 2.
@@
    As a result, the mean square error $\mathsf{E}(\|{\mathbf{s}}-\tilde{{\mathbf{s}}}\|_{2}^{2})$ is lower bounded by $\mathsf{Tr}({\mathbf{I}}^{-1})$, where $\mathsf{Tr}(\cdot)$ denotes the trace operator. In Section V, we numerically show that our proposed algorithms can approach the above CRLB.
    IV-D2 Computational Complexity Analysis
@@
    V Simulation Results
    V-A Experimental Settings
@@
    The default simulation setups for atomic MIMO receivers are as follows unless specified otherwise. The number of atomic antennas is $N=36$, while the number of single-antenna users is $K=3$. The 4-QAM and 16-QAM modulators are considered. The Rydberg energy levels $52D_{5/2}$ and $53P_{3/2}$ are adopted for detecting the sub-6G signals of frequency $\omega_{eg}=2\pi\times 5$ GHz. According to the analytical result in [12], the length of the dipole matrix element, i.e., ${\mathbf{r}}_{eg}^{H}\bm{\epsilon}_{nkl}$ in (14), is of the order $(0.5\sim 1.5)n^{2}a_{0}$, where $a_{0}=5.292\times 10^{-11}$ m specifies the Bohr radius and $n$ is 52 in our configuration. Taking into account the randomness of the polarization direction, we use a Gaussian distribution $\mathcal{CN}\left(0,n^{4}a_{0}^{2}\right)$ to generate ${\mathbf{r}}_{eg}^{H}\bm{\epsilon}_{nkl}$. In addition, the channel coefficients are generated using the standard 3GPP TR 38.901 model, whose key parameters are given in Table I. The received SNR is defined as
    ${\rm SNR}=\frac{\mathsf{E}(|{\mathbf{a}}_{n}^{H}{\mathbf{s}}|^{2})}{\mathsf{E}(|w_{n}|^{2})}=\frac{\sum_{k=1}^{K}\mathsf{E}(|a_{nk}s_{k}|^{2})}{\sigma^{2}}.$ (32)
@@
    • ZF with known phase: Assume the ideal case where the true phase of ${\mathbf{y}}$ is known in advance. Then the zero-forcing (ZF) detector is employed to recover the data signal, ${\mathbf{s}}$.
@@
    V-B Evaluation of NMSE Performance
@@
    V-C Evaluation of BER Performance
    To evaluate the BER performance, the constellation demapping step is introduced to project the recovered signals $\tilde{s}_{k}$ for each user to the nearest constellation point $\hat{s}_{k}$ and then to 0-1 bits. The curves of BER versus SNR are plotted in Fig. 7. The RSR is set as $12$ dB. Fig. 7(a) adopts a small-scale configuration: $N\times K=36\times 3$ with a 4-QAM modulator, while Fig. (b) employs a large-scale configuration: $N\times K=100\times 6$ with a 16-QAM modulator. It can be seen from Fig. 7(a) that the biased GS and the EM-GS algorithms exhibit a remarkable BER reduction compared with CM-ZF and can perform similarly as the exhaustive search method with much lower complexity. Furthermore, the EM-GS algorithm consistently outperforms the biased GS algorithm due to the high-pass filter $R(\cdot)$ (see their comparison in Section IV-C). For the large-scale configuration depicted in Fig. 7(b), the computation of exhaustive search method is prohibitive, so it is excluded from this comparison. We can observe that the SNR gap between the EM-GS algorithm and the ZF method with known phase for realizing the same BER level is between $3\sim 4$ dB. This fact indicates that despite the lack of phase information, our EM-GS based detector can attain a similar BER trend as the ZF detector in the ideal case, demonstrating the effectiveness of multi-user atomic-MIMO detection.
@@ VI Conclusions
    $i\hbar\left(\begin{array}[]{c}\dot{\alpha}_{e,n}\\ \dot{\alpha}_{g,n}\end{array}\right)=\left(\begin{array}[]{cc}\hbar\omega_{e}&\xi_{n}\\ \xi_{n}^{*}&\hbar\omega_{g}\end{array}\right)\left(\begin{array}[]{c}{\alpha}_{e,n}\\ {\alpha}_{g,n}\end{array}\right).$ (40)
    To solve this time-varying differential equation, the first step is to switch (40) into the interaction picture of signal frequency. Specifically, let $\widetilde{\alpha}_{e,n}=e^{i(\omega+\omega_{g})t}\alpha_{e,n}$ and $\widetilde{\alpha}_{g,n}=e^{i\omega_{g}t}\alpha_{g,n}$. By applying the relationships in (40), the differential equation of $\widetilde{\alpha}_{e,n}$ is derived as
@@
    and the differential equation of $\widetilde{\alpha}_{g,n}(t)$ is similarly derived as
@@
    where $\xi_{n}^{-}=\frac{1}{2}\sum_{k}\sum_{l}{{\mathbf{r}}}_{eg}^{T}\bm{\epsilon}_{nkl}q\sqrt{P_{k}}\rho_{nkl}|s_{k}|e^{-j(\varphi_{nkl}+\gamma_{k})}$ is a time-independent term while the time-dependent term $\xi_{n}^{+}=\frac{1}{2}\sum_{k}\sum_{l}{{\mathbf{r}}}_{eg}^{T}\bm{\epsilon}_{nkl}q\sqrt{P_{k}}\rho_{nkl}|s_{k}|e^{j(\varphi_{nkl}+\gamma_{k})}e^{j2\omega t}$ oscillates with frequency $2\omega$. Due to the rapid oscillation of $\xi_{n}^{+}$, the RWA states that the time integrals of $\xi_{n}^{+}\widetilde{\alpha}_{g,n}$ and ${\xi_{n}^{+}}^{*}\widetilde{\alpha}_{e,n}$ quickly averages to zero, and thus can be neglected. Thereafter, the equations (-A) and (42) are approximated as
@@
    where $\Phi_{n}=\arctan(\frac{2|\xi_{n}^{-}|}{\hbar\delta})$ and $\Psi_{n}=\angle(\xi_{n})$. Thereafter, the evolution of $(\widetilde{\alpha}_{e,n},\widetilde{\alpha}_{g,n})$ driven by (50) is expressed as
@@
    Here, the coefficients $c_{n}^{\pm}$ are determined by the initial condition. Given the initial states $\widetilde{\alpha}_{e,n}(0)=\alpha_{e,n}(0)=0$ and $\widetilde{\alpha}_{g,n}(0)={\alpha}_{g,n}(0)=1$, associated with the initial derivative $\dot{\widetilde{\alpha}}_{e,n}(0)=-i\delta\widetilde{\alpha}_{e,n}(0)-i\frac{\xi_{n}^{-}}{\hbar}\alpha_{g,n}(0)=-i\frac{\xi_{n}^{-}}{\hbar}$, the probability density is obtained as
@@
    Substituting (26) and (27) into (IV-B), the surrogate function can be expressed as

### v3

@@ III-C Noise Model
IV Atomic-MIMO Signal Detection
In this section, we first formulate the signal detection problems of atomic MIMO receivers based on the LS and ML principles. Then, two atomic-MIMO detection algorithms are proposed to solve the formulated problems. Their performance and complexity are subsequently analyzed.
IV-A Atomic-MIMO Signal Detection Problems
The objective of atomic-MIMO signal detection is to estimate the multi-user data vector, ${\mathbf{s}}$, from the received signal, ${\mathbf{z}}$, in (23). Two problem formulations are provided below.
IV-A1 Least-Square Criterion
@@
IV-A2 Maximum-Likelihood Criterion
@@
Although problems (24) and (26) can be optimally solved by an exhaustive search, the complexity is prohibitive when the users are many and/or the constellation size is large. To this end, two low-complexity algorithms will be designed subsequently.
@@
IV-B Atomic-MIMO Signal Detection Based on LS Criterion
One issue in the implementation of biased GS algorithm is its sensitivity to the choice of initial point. Without a proper initial estimate, Algorithm 1 may converge to a shallow local optimum. To address this issue, the computationally efficient spectral method has been developed to yield good initialization [33]. Specifically, to adapt the spectral method into our model, we first construct an augmented observation matrix $\bar{{\mathbf{A}}}=[{\mathbf{A}}^{H},{\mathbf{b}}]^{H}$ as presented in step 1. This matrix allows us to treat the user signal and reference signal together: $\bar{{\mathbf{s}}}=[{\mathbf{s}}^{H},1]^{H}$. Thereafter the observed amplitude vector can be rewritten in a reference-free form: ${\mathbf{z}}=|{\mathbf{A}}^{H}{\mathbf{s}}+{\mathbf{b}}+{\mathbf{w}}|=|\bar{{\mathbf{A}}}^{H}\bar{{\mathbf{s}}}+{\mathbf{w}}|$, whose initialization can be performed using the typical spectral method. To elaborate, define the weighted covariance matrix as ${\mathbf{M}}=\sum_{n=1}^{N}z_{n}\bar{{\mathbf{a}}}_{n}\bar{{\mathbf{a}}}_{n}^{H}$, where $\bar{{\mathbf{a}}}_{n}$ denotes the $n$-th column of $\bar{{\mathbf{A}}}$. The direction of the initial value $\bar{{\mathbf{s}}}^{0}$ of $\bar{{\mathbf{s}}}$ is estimated as the principal eigenvector ${\mathbf{v}}$ of ${\mathbf{M}}$. In addition, the magnitude $\bar{r}$ of $\bar{{\mathbf{s}}}^{0}$ is obtained by approaching $|{\mathbf{A}}^{H}\bar{{\mathbf{s}}}^{0}|=|{\mathbf{A}}^{H}\bar{r}{\mathbf{v}}|$ to ${\mathbf{z}}$:
@@ IV-B Atomic-MIMO Signal Detection Based on LS Criterion
IV-C Atomic-MIMO Signal Detection Based on ML Criterion
@@
where ${\mathbf{s}}^{t-1}$ denotes the previous estimate of ${\mathbf{s}}$. The EM algorithm proceeds iteratively by two steps to maximize the loglikelihood. They are the Expectation step (E-step) to derive the surrogate function $Q({\mathbf{s}}|{\mathbf{s}}^{t-1})$ by calculating the expectation in (IV-C), and the Maximization step (M-step) to update ${\mathbf{s}}^{t}$ by maximizing $Q({\mathbf{s}}|{\mathbf{s}}^{t-1})$, i.e., ${\mathbf{s}}^{t}=\arg\max_{{\mathbf{s}}}Q({\mathbf{s}}|{\mathbf{s}}^{t-1})$. These two steps are elaborated below.
  • E-step: In (IV-C), the likelihood function $p(z_{n},\theta_{n};{\mathbf{s}})$ stands for the ML estimator to ${\mathbf{s}}$ when both the amplitude $z_{n}$ and phase $\theta_{n}$ are accessible. Clearly, $p(z_{n},\theta_{n};{\mathbf{s}})$ is a Gaussian distribution
@@
    Moreover, to avoid convergence to a shallow local optimum, the spectral method based on the augmented matrix $\bar{{\mathbf{A}}}$ as used in Section IV-B is carried out to obtain an initial estimate ${\mathbf{s}}^{0}$. In the end, after performing the initialization and EM-based iteration, the projection operation $\hat{s}_{k}=\arg\min_{s\in\mathcal{S}}\|\tilde{s}_{k}-s\|$ is applied to complete the constellation demapping.
  • IV-D Comparison between biased GS and EM-GS algorithms
@@
    IV-E Performance Analysis
    IV-E1 CRLB Analysis
@@
    where ${\mathbf{I}}$ denotes the Fisher information matrix, the $(p,q)$-th entry of which is determined by $I_{pq}=-\sum_{n=1}^{N}\mathsf{E}\left(\frac{\partial^{2}}{\partial s_{p}^{*}\partial s_{q}}\log p(z_{n};{\mathbf{s}})\right)$. An explicit expression of ${\mathbf{I}}$ is provided by the following lemma.
@@
    As a result, the mean square error $\mathsf{E}(\|{\mathbf{s}}-\tilde{{\mathbf{s}}}\|_{2}^{2})$ is lower bounded by $\mathsf{Tr}({\mathbf{I}}^{-1})$, where $\mathsf{Tr}(\cdot)$ denotes the trace operator. In Section VI, we numerically show that our proposed algorithms can approach the above CRLB.
    IV-E2 Computational Complexity Analysis
@@
    V Extension to Channel Estimation
    In previous discussions, we assume that the channel state information (CSI) is perfectly known to the receiver. In case the perfect CSI is not available, we can estimate the CSI, $\mathbf{A}$, by sending pilot signals to the receiver.
    Consider the received signal at the $n$-th antenna: $z_{n}=|\mathbf{a}_{n}^{H}\mathbf{s}+b_{n}+w_{n}|=|\mathbf{s}^{H}\mathbf{a}_{n}+b_{n}^{*}+w_{n}^{*}|$, where $\mathbf{a}_{n}$ is the channel vector to be estimated. In practice, the LO is deployed at a fixed position near the atomic receiver. Thus, the coefficient $b_{n}$ almost keeps unchanged and can be obtained in advance by measuring the physical locations of LO and receiver. To estimate the channel coefficients $\mathbf{a}_{n}$, we let the $K$ users send known pilot signals during $P$ time slots to the receiver, which are denoted as $\mathbf{s}_{p}\in\mathbb{C}^{K\times 1},p\in\{1,\cdots,P\}$. The overall received signal $\mathbf{z}_{n}\in\mathbb{R}^{P\times 1}$ is thus expressed as
    $\mathbf{z}_{n}$ $=|[\mathbf{s}_{1},\cdots,\mathbf{s}_{P}]^{H}\mathbf{a}_{n}+[b_{n},\cdots,b_{n}]^{H}+[w_{n,1},\cdots,w_{n,P}]^{H}|$
    $=|\mathbf{S}^{H}\mathbf{a}_{n}+\mathbf{b}_{n}+\mathbf{w}_{n}|,$ (37)
    where $\mathbf{S}=[\mathbf{s}_{1},\dots,\mathbf{s}_{P}]\in\mathbb{C}^{K\times P}$ refers to the known pilot matrix, $\mathbf{b}_{n}=[b_{n},\cdots,b_{n}]^{H}\in\mathbb{C}^{P\times 1}$ the reference signal, and $\mathbf{w}_{n}=[w_{n,1},\cdots,w_{n,P}]^{H}\in\mathbb{C}^{P\times 1}$ the noise. It is clear that the channel estimation model in (V) for recovering ${\mathbf{a}}_{n}$ from ${\mathbf{z}}_{n}$ has the same mathematical structure as the signal detection model ${\mathbf{z}}=|{\mathbf{A}}^{H}{\mathbf{s}}+{\mathbf{b}}+{\mathbf{w}}|$ for detecting ${\mathbf{s}}$ from ${\mathbf{z}}$. Thereafter, both the biased GS and EM-GS algorithm can be employed to estimate the channel coefficients ${\mathbf{a}}_{n}$. In the end, this process can be simultaneously carried out on all receive antennas to recover the whole channel matrix ${\mathbf{A}}=[{\mathbf{a}}_{1},{\mathbf{a}}_{2},\cdots,{\mathbf{a}}_{N}]$. This completes the channel estimation task.
    VI Simulation Results
    VI-A Experimental Settings
@@
    The default simulation setups for atomic MIMO receivers are as follows unless specified otherwise. The number of atomic antennas is $N=36$, while the number of single-antenna users is $K=3$. The 4-QAM and 16-QAM modulators are considered. The Rydberg energy levels $52D_{5/2}$ and $53P_{3/2}$ are adopted for detecting the sub-6G signals of frequency $\omega_{eg}=2\pi\times 5$ GHz. Utilizing the Python package [36], the transition dipole moment $\bm{\mu}_{eg}$ over the states $52D_{5/2}$ and $53P_{3/2}$ is calculated as $[0,1785.916\>qa_{0},0]^{T}$, where $a_{0}=5.292\times 10^{-11}$ m specifies the Bohr radius. Taking into account the randomness of the polarization direction, vectors $\bm{\epsilon}_{nkl}$ and $\bm{\epsilon}_{b,n}$ are randomly sampled from unit circles perpendicular to their incident angles. In addition, the channel coefficients are generated using the standard 3GPP TR 38.901 model, whose key parameters are given in Table I. The received SNR is defined as
    ${\rm SNR}=\frac{\mathsf{E}(|{\mathbf{a}}_{n}^{H}{\mathbf{s}}|^{2})}{\mathsf{E}(|w_{n}|^{2})}.$ (38)
@@
    • ZF with known phase: Assume the ideal case where the true phase of ${\mathbf{y}}$ is known in advance. Then the zero-forcing (ZF) detector is employed to recover the QAM symbol, ${\mathbf{s}}$.
@@
    VI-B Evaluation of NMSE Performance
@@
    VI-C Evaluation of BER Performance
    To evaluate the BER performance, the constellation demapping step is introduced to project the recovered signals $\tilde{s}_{k}$ for each user to the nearest constellation point $\hat{s}_{k}$ and then to 0-1 bits. The curves of BER versus SNR are plotted in Fig. 7. The RSR is set as $12$ dB. Fig. 7(a) adopts a small-scale configuration: $N\times K=36\times 3$ with a 4-QAM modulator, while Fig. 7(b) employs a large-scale configuration: $N\times K=100\times 6$ with a 16-QAM modulator. It can be seen from Fig. 7(a) that the biased GS and the EM-GS algorithms exhibit a remarkable BER reduction compared with CM-ZF and can perform similarly as the exhaustive search method with much lower complexity. Furthermore, the EM-GS algorithm consistently outperforms the biased GS algorithm due to the high-pass filter $R(\cdot)$ (see their comparison in Section IV-D). For the large-scale configuration depicted in Fig. 7(b), the computation of exhaustive search method is prohibitive, so it is excluded from this comparison. We can observe that the SNR gap between the EM-GS algorithm and the ZF method with known phase for realizing the same BER level is between $3\sim 4$ dB. This fact indicates that despite the lack of phase information, our EM-GS based detector can attain a similar BER trend as the ZF detector in the ideal case, demonstrating the effectiveness of multi-user atomic-MIMO detection.
@@ VII Conclusions
    $i\hbar\left(\begin{array}[]{c}\dot{\alpha}_{e,n}\\ \dot{\alpha}_{g,n}\end{array}\right)=\left(\begin{array}[]{cc}\hbar\omega_{e}&\xi_{n}\\ \xi_{n}^{*}&\hbar\omega_{g}\end{array}\right)\left(\begin{array}[]{c}{\alpha}_{e,n}\\ {\alpha}_{g,n}\end{array}\right),$ (46)
    where $\dot{\alpha}$ denotes the time derivative of $\alpha$. To solve this time-varying differential equation, the first step is to switch (46) into the interaction picture of signal frequency. Specifically, let $\widetilde{\alpha}_{e,n}=e^{i(\omega+\omega_{g})t}\alpha_{e,n}$ and $\widetilde{\alpha}_{g,n}=e^{i\omega_{g}t}\alpha_{g,n}$. By applying the relationships in (46), the differential equation of $\widetilde{\alpha}_{e,n}$ is derived as
@@
    and the differential equation of $\widetilde{\alpha}_{g,n}$ is similarly derived as
@@
    where $\xi_{n}^{-}=\frac{1}{2}\sum_{k}\sum_{l}{\bm{\mu}}_{eg}^{T}\bm{\epsilon}_{nkl}\sqrt{P_{k}}\rho_{nkl}|s_{k}|e^{-j(\varphi_{nkl}+\gamma_{k})}$ is a time-independent term while the time-dependent term $\xi_{n}^{+}=\frac{1}{2}\sum_{k}\sum_{l}{\bm{\mu}}_{eg}^{T}\bm{\epsilon}_{nkl}\sqrt{P_{k}}\rho_{nkl}|s_{k}|e^{j(\varphi_{nkl}+\gamma_{k})}e^{j2\omega t}$ oscillates with frequency $2\omega$. Due to the rapid oscillation of $\xi_{n}^{+}$, the RWA states that the time integrals of $\xi_{n}^{+}\widetilde{\alpha}_{g,n}$ and ${\xi_{n}^{+}}^{*}\widetilde{\alpha}_{e,n}$ quickly averages to zero, and thus can be neglected. Thereafter, the equations (-A) and (48) are approximated as
@@
    where $\Phi_{n}=\arctan(\frac{2|\xi_{n}^{-}|}{\hbar\delta})$ and $\Psi_{n}=\angle(\xi_{n}^{-})$. Thereafter, the evolution of $(\widetilde{\alpha}_{e,n},\widetilde{\alpha}_{g,n})$ driven by (56) is expressed as
@@
    Here, the coefficients $c_{n}^{\pm}$ are determined by the initial condition. Given the initial states $\widetilde{\alpha}_{e,n}(0)=\alpha_{e,n}(0)=0$ and $\widetilde{\alpha}_{g,n}(0)={\alpha}_{g,n}(0)=1$, associated with the initial derivative $\dot{\widetilde{\alpha}}_{e,n}(0)=-i\delta\widetilde{\alpha}_{e,n}(0)-i\frac{\xi_{n}^{-}}{\hbar}\widetilde{\alpha}_{g,n}(0)=-i\frac{\xi_{n}^{-}}{\hbar}$, the probability density is obtained as
@@
    Substituting (31) and (32) into (IV-C), the surrogate function can be expressed as